Cash flo v to o s. Use tho dena f the fo ng finere al staterments from ownars (change in common stock and paid-in-aurplus]. The dstbutod oamings t te popu wndo . The con per y pad merest expe ee o 817,400 fr 2017 and had an o emi tox a e c 40% fr 2017. Find the es fo to o ers fr 2017 by perta end b tal with the pats being d de da ped d nenes es n bom ving Data Table Round to the noarost dalar.) Yoar End 350,000 ococds coid Fixod costs Soling. general, and administrativo experses Ceprecietion $28.00(j S6,100 Sheet 12/31/2016 LIABILITIES 18,1DD Nctos payabla S13, $19,1 $189 2780Acounts payablo $48,200 Long-tam debt $367,900OWNERS EQUITY 142,200Rctainod earnings 81,800 Common stook gibla assots $131 12/3112017 525,800 Nctes peyable $18,000 Accounts payable 553,200 Long-term debt $182,1 448,100OWNERS EQUITY arnings sses 581.900Common stock 5182 Enter your answer in the arswer box and then click Check Answer Print Done Check Answer
Answer #1
1) The first step is to complete the balance sheets of the two years by
finding out the missing figures.
2016 2017
Cash 16100 25800
Accounts receivable 27800 18900
Inventories 48200 53200
Total current assets 92100 97900
Fixed assets 367800 448100
Accumulated depreciation -142200 -188300
Net Fixed Assets 225600 259800
Intangible assets 81800 81800
Total assets 399500 439500
Notes payable 13900 11800
Accounts payable 19100 24200
Current liabilities 33000 36000
Long term debt 189900 162100
Retained earnings (plug in figure) 44800 59400
Common stock 131800 182000
Total liabilities and equity 399500 439500
OTHER WORK SHEETS AND CALCULATIONS:
INCOME STATEMENT FOR 2017
Sales revenue 350000
COGS 142000
Fixed costs 43200
Selling, general and administrative expenses 28000
Depreciation 46100
EBIT 90700
Interest 17400
EBT 73300
Tax at 40% 29320
NI 43980
Dividend paid during 2017 = Beginning retained earnings+Net income-Ending retained earnings = 44800+43980-59400 = 29380
CASH FLOWS:
Cash flows from assets = Cash flow to creditors+Cash flow stock holders.
1) Cash flow from assets:
Cash flow from assets = Operating cash flow-Net capital spending-Change in NWC
Operating cash flow = NI+Depreciation+Interest expense = 43980+46100+17400 = 107480
Net capital spending = Ending net fixed assets-Beginning net fixed assets+Depreciation expense = 259800-225600+46100 = 80300
Change in NWC = Ending NWC-Beginning NWC = (97900-36000)-(92100-33000)= 2800
Cash flow from assets = 107480-80300-2800 = 24380
2) Cash flow to creditors = Interest expense-(Ending LT loan balance-Beginning LT loan balance) = 17400-(162100-189900)= 45200
3) Cash flow to owners = Dividends paid-(Ending common stock-Beginning common stock) = 29380-(182000-131800) = -20820
4) Cash flow to creditors+Cash flow to owners = 45200-20820 = 24380
5) CASH FLOW FROM ASSETS = CASH FLOW TO CREDITORS+CASH FLOW TO OWNERS